OpenStudy (anonymous):

is there a fast way to solve this problem? tickets to a benefit concert were sold at $80 and $50 a piece respectively. the total receipts being $6200. if twice as many 80ticket and half as many 50ticket had been sold, the receipts would have been $7900. how many $80 tickets were sold?

OpenStudy (amistre64):

80a+50b = 6200 2(80a) + (50b)/2 = 7900 80a + 50b = 6200 ; *-2 160a + 25b = 7900 -160a -100b = -12400 160a + 25b = 7900 ---------------------- -75b = -4500 b@50 = 4500/74 ; if i did it right

OpenStudy (amistre64):

i spose the smarter step would have been to eliminate the b and solve for a :)

OpenStudy (amistre64):

80a + 50b = 6200 160a + 25b = 7900 ; *-2 80a +50b = 6200 -320a -50b = -15800 --------------------- -240a = - 9600 a = 9600/240 perhaps?

OpenStudy (amistre64):

dbl chk 40 @ $80

OpenStudy (anonymous):

please help me with this too! http://openstudy.com/users/phoebeco#/users/phoebeco/updates/4e2ee50c0b8b63299a425450